# LLM-ARK: Knowledge Graph Reasoning Using Large Language Models via Deep Reinforcement Learning

### 18 Dec 2023 v1

@@ 3.3.3 Training
1. Optimizer We propose leveraging the experiences accumulated by LLM-ARK to execute KG reasoning. More formally, for the above policy network ($\pi_{\theta}$), we want to find the parameter $\theta$ that maximizes the reward.
$\begin{split}J(\theta)=\mathbb{E}_{\left(e_{s},\mathcal{P},e_{g}\right)\sim D}\mathbb{E}_{A_{1},\ldots,A_{T-1}\sim\pi_{\theta}}\\ \left[R\left(S_{T}\right)\mid S_{1}=\left(FTE_{1}\right)\right],\end{split}$ (5)
@@
$\mathbf{V}$ $=\mathbf{W}_{\mathbf{3}}\left(\operatorname{tanh}\left(\mathbf{W}_{\mathbf{2}}\operatorname{tanh}\left(\mathbf{W}_{\mathbf{1}}\left(s_{t}\right)\right)\right)\right)$ (6)
@@
2. Replay Buffer The replay buffer stores the triplets $rb=(c^{e},s,l,a,s^{{}^{\prime}},done)$ of the reflection prompt, indicating the current entity, the current state, logits, the selected action, the next state, and whether the episode has ended. The reason for recording the current entity is that we need to get all exit paths of the current entity for further fusion 4.
@@ 4.1 Datasets
OpenDialKG is a publicly available parallel corpus, consisting of 91,000 dialogues, each of which is supplemented with paths that link knowledge graph (KG) entities and their relations. The aim of the corpus is to render human dialogues’ implicit reasoning processes as explicit computer operations within the KG. The dataset is partitioned into a 70% training set, 15% validation set, and 15% test set, following the methodology described in Moon et al. (2019).
@@ 4.2 Baselines
We compared our results with these baseline models: Tri-LSTM, Seq2Seq, Seq2Path, DialKG Walker(Young et al., 2018; Moon et al., 2019), DiffKG, AttnFlow, AttnIO(Jung et al., 2020) and HiTKG(Ni et al., 2022). HiTKG is a hierarchical transformer-based tool that uses diverse inputs to forecast knowledge graph paths. Our team is highly influenced by this research and has chosen HiTKG as a strong baseline. To evaluate the performance of the state-of-the-art LLMs on the knowledge graph inference task, we designed three prompts methods: GPT4-Standard, GPT4-Normal and GPT4-OPA. The difference between GPT4-Standard, GPT4-Normal is that GPT4-Normal has more awareness of dialog context and path history, while GPT4-OPA has more awareness of 2-hop exit path subgraphs compared to the former two. See appendix A.3 for full prompts.
@@ 4.3 Implement Details
The training was conducted on A40. Informed by prior research from Jung et al. (2020); Ni et al. (2022), we pre-trained the knowledge graph using TransE(Fan et al., 2014) based on this GitHub repository [^3] https://github.com/thunlp/OpenKE. The objective was to unearth and explore entity relationships, expand the knowledge network for connection prediction, and enable diverse reward function design. To reduce GPU memory usage and increase the pace of training, all experiments - excluding LLaMA7B-ARK-FP32 were carried out with bfloat16 computational data type. We selected LLaMA-7B and LLaMA-13B, and since all true paths in OpenDialKG are at most 2 hops, we set the maximum path length to $T=2$. We included ”Equal” to ensure that the model stops automatically after the second hop. To ensure fairness, we randomly shuffled the exit paths of the knowledge graph. We set max patience to 20, meaning that training is terminated if there is no boost for 20 rewards on the validation set. Further information on the hyperparameters is available in the Appendix A.
@@ 4.5 Comparative Experiments
For the KG reasoning task, we assessed path recall at different K values (1, 3, 5, 10, 25) and target entity recall at position K (1, 3, 5, 10, 25). Our findings, which are presented in Table 1, demonstrate that our proposed LLM-ARK performs better than all benchmarked baselines in target@k metrics, surpassing the GPT-4. The performance gain is signifcant, especially in recalls with taget@k: there is a 17.64% relative improvement in target@1, 17.44% in target@5 and 16.56% in target@10. Unfortunately, our model’s path@k evaluation matrix socores do not outperform the state-of-the-art (SOTA) model because we trained using only the target arrival reward function, but we are very extensible and there is potential for improvement.
The GPT-4-Standard and GPT-4-Normal methods are deficient in path awareness. Nonetheless, our models exhibit improved outcomes with the addition of path awareness. Furthermore, GPT-4 rely solely on Zero Shot decoding for output path decoding. In contrast, our model implements PA-MLP to improve the environment-aware capabilities of LLMs. Additionally, due to the substantial comprehension of instructions and comprehensive information encoding features of the large model, we can modify cue words, eliminating the need for the complex multi-input aggregation technique utilized in HiTKG. Viewing large models as intelligent agents that explore a knowledge graph to acquire experience can reap benefits from the positive-negative feedback optimization mechanism of the Reinforcement Learning Policy Supervisor Algorithm. This methodology strengthens the training of our model to perform flexible reasoning on KGs in multi-step scenarios, outperforming not only GPT-4 but also smaller models. The efficacy of our approach is corroborated by our model’s superior performance.
@@
LLaMA7B-ARK(proposed) 22.37 32.79 38.92 49.45 60.18 48.75 63.73 72.15 84.12 90.50
LLaMA13B-ARK 22.53 31.40 37.29 46.67 55.45 48.44 61.43 69.52 81.00 88.71
LLaMA7B-ARK-UI 17.27 26.06 32.04 41.53 50.27 46.64 62.22 71.28 81.88 89.49
LLaMA7B-ARK-FP32 21.36 28.34 33.30 41.30 50.62 42.38 52.91 61.64 74.89 87.25
LLaMA7B-ARK-WT 16.13 21.42 27.40 36.79 47.41 32.88 46.24 55.69 71.12 84.32
LLaMA7B-ARK-WP 1.05 2.31 3.26 4.62 5.66 41.54 63.21 75.07 87.22 96.47
@@
4.6 Analysis Experiment
As shown in Table 2, LLM-ARK was benchmarked against multiple ablation models on the OpenDialKG dataset. To evaluate the impact of instructions on model performance, we trained the LLaMA7B-ARK-UI model without instruction. The results of this model are the closest to those of LLaMA7B-ARK, indicating that the presence or absence of commands has an effect on the model’s results, but not a serious one. To conserve computational resources, we use the bfloat16 computational type for all experiments except for LLaMA7B-ARK-FP32. This is due to the fact that LLMs require more GPU memory space, necessitating a reduction in the batchsize and a slower training speed. To our surprise, the Float32 computational type did not achieve higher experimental results. To determine the impact of model parameter size on training outcomes, experiments were conducted utilizing the LLaMA7B and LLaMA13B models. The model LLaMA7B-ARK-FP32 with a greater number of parameters only achieve higher scores for path@1. The KG reasoning task does not demand the model to execute intricate commands, thus a 7B parameter-sized model can effectively fit the KG reasoning dataset. LLaMA7B-ARK-WT that was trained by randomly initializing relation and entity embedding for the LLaMA7B-ARK-WT model is less effective than LLaMA7B-ARK. We conducted ablation experiments LLaMA7B-ARK-WP and found that the performance of our export environment-aware sub-module PA-MLP decreases substantially if we do not consider the exit paths. The score for path@1 is only 1.05%, which is 21.32% lower than LLaMA7B-ARK. These ablation experiments and results demonstrate the contribution and necessity of our subcomponents to the model.
4.7 Case Study
@@ 5 Conclusion
Our study explores the potential of LLMs in KG reasoning tasks through an examination of GPT-4 (OpenAI, 2023) and an understanding of the basic capabilities of LLMs. Our model, by conducting flexible and precise learning with rich environmental information, enabled flexible and accurate multilevel inference, attaining superior results compared to both GPT-4, and smaller models. Ultimately, our experimental results affirmed that misalignment between model knowledge and environmental data can significantly undermine model performance. Experiments also affirmed the critical role of considering all aspects in the model optimization process. We hope our research provides a valuable contribution to the field and inspires future work in this area.
Acknowledgments
We thank Lida Shi, Anqi Liu and members of the HAI Deep Learning group for valuable discussions and comments.
@@ A.2 Tricks
In the original paper of PPO, no implementation details and techniques are mentioned other than the use of GAE to compute the dominance function. Referring to this repository [^4] https://github.com/Lizhi-sjtu/DRL-code-pytorch, we employ several optimization tricks.In the actual code implementation, to encourage the diversity of paths sampled by the strategy during training, we added an entropy regularization term to our loss function, scaled by the constant ($\beta$). We used the operation of normalization of advantage proposed in the paper [Tucker et al., 2018]. Learning rate decay can enhance the smoothness in the late training stage to some extent and improve the training effect. Here we use the linear decay of learning rate, with the number of training steps learning rate from the initial value of a linear decline to 0. Gradient clipping is a trick introduced to prevent the gradient from exploding during the training process, which also serves to stabilize the training process. Orthogonal Initialization is a neural network initialization method proposed to prevent problems such as gradient vanishing and gradient explosion at the beginning of training. Referring to the MAPPO [Yu et al., 2022], the Adam optimizer individually sets eps=1e-5, and this particular setting can improve the training performance of the algorithm to some extent.
A.3 GPT4 Prompts
For GPT4-OPA prompts, since max length is 2, we need to recursively get all exit paths at the next level of all exit paths of the current entity, most of which are omitted due to the large number of KG subgraph triples of exit paths.
@@ A.3 GPT4 Prompts
### Environment Utterance: What do you think about the Washinton Redskins? Are you a fan? Current Entity: Washington Redskins
@@
Path History: [] Current Entity: Washington Redskins
@@
Path History: [] Current Entity: Washington Redskins
@@ A.4 HyperParameters
train times 8
batchsize 4096
@@
train batch size 4
test batch size 48

### 04 Feb 2024 v2

@@ 3.3.3 Training
Optimizer Our model is optimized by utilizing the experience accumulated by agent during KG reasoning. More formally, for the above policy network ($\pi_{\theta}$), we want to find the parameter $\theta$ that maximizes the reward.
$\begin{split}J(\theta)=\mathbb{E}_{\left(e_{s},\mathcal{P},e_{g}\right)\sim D}\mathbb{E}_{A_{1},\ldots,A_{T-1}\sim\pi_{\theta}}\\ \left[R\left(S_{t}\right)\mid S_{1}=\left(s_{1}\right)\right],\end{split}$ (5)
@@
$\mathbf{V}$ $=\mathbf{W}_{\mathbf{3}}\left(\operatorname{tanh}\left(\mathbf{W}_{\mathbf{2}}\left(\operatorname{tanh}\left(\mathbf{W}_{\mathbf{1}}\left(s_{t}\right)\right)\right)\right)\right)$ (6)
@@
Replay Buffer The replay buffer stores the triplets $rb=(v_{c},s,logit,a,s^{{}^{\prime}},done)$ of the reflection prompt, indicating the current entity, the current state, logits, the selected action, the next state, and whether the episode has ended. The reason for recording the current entity is that we need to get all exit paths of the current entity for further fusion 4.
@@ 4.1 Datasets
OpenDialKG is a publicly available parallel corpus of conversations and Knowledge Graphs consisting of 91,000 conversations, each supplemented by paths connecting Knowledge Graph entities and their relationships. The purpose of the corpus is to present the implicit reasoning processes of human dialog as explicit computer operations on the Knowledge Graph. Following previous work described in Moon et al. (2019), we split this dataset into a 70% training set, a 15% validation set, and a 15% test set.
@@ 4.2 Baselines
We compared our results with these baseline models: Tri-LSTM, Seq2Seq, Seq2Path, DialKG Walker(Young et al., 2018; Moon et al., 2019), DiffKG, AttnFlow, AttnIO(Jung et al., 2020) and HiTKG(Ni et al., 2022). HiTKG is a hierarchical transformer-based tool that uses diverse inputs to predict KG paths. Our team chose HiTKG as a strong baseline. To evaluate the performance of the state-of-the-art LLMs on the KG inference task, we designed three prompt methods: GPT4-Standard, GPT4-Normal and GPT4-OPA. The difference between GPT4-Standard, GPT4-Normal is that GPT4-Normal has more awareness of dialog context and path history, while GPT4-OPA has more awareness of 2-hop exit path subgraphs compared to the former two. See appendix A.4 for full prompt.
@@ 4.3 Implement Details
The training was conducted on A40. Informed by prior research from Jung et al. (2020); Ni et al. (2022), we pre-trained the knowledge graph using TransE (Fan et al., 2014) based on this GitHub repository [^2] https://github.com/thunlp/OpenKE. The objective was to unearth and explore entity relationships, expand the knowledge graph for connection prediction, and enable diverse reward function design. To facilitate reproducibility, we adopt an open-source LLM, i.e., LLaMA-2-7B(Touvron et al., 2023b). To reduce GPU memory usage and increase the pace of training, all experiments - excluding LLaMA-2-7B-ARK-FP32 were carried out with BFLOAT16(Kalamkar et al., 2019) half-precision format. Since all true paths in OpenDialKG are at most 2 hops, we set the maximum path length to $t=2$. We included ”Equal” to ensure that the model stops automatically after the second hop. To ensure fairness, we randomly shuffled the exit paths of the knowledge graph. We set max patience to 5, meaning that training is terminated if there is no boost for 5 rewards on the validation set. Further information on the hyperparameters is available in the Appendix 8.
@@ 4.5 Comparative Experiments
For the KG reasoning task, we assessed path recall at different K values (1, 3, 5, 10, 25) and target entity recall at position K (1, 3, 5, 10, 25). As presented in Table 1, the result demonstrate that our proposed model LLaMA-2-7b-ARK performs better than all benchmarked baselines in target@1, 5, 10, 25 metrics. The performance gain is signifcant, especially in recalls with taget@1, 10: there is a 5.28% relative improvement in target@1 and 9.59% in target@10. Unfortunately, our model’s path@k evaluation matrix socores do not outperform the current state-of-the-art (SOTA) model HiTKG because we trained using only the target arrival reward function, but we are very extensible and there is potential for improvement. As described in this paper, we evaluate the performance of GPT-4 in performing dialog inference using manual prompt constructed with different environmental information. Therefore, we also report the performance of GPT-4 with different prompts on the same dataset.
At the decoding stage of AttnIO, AttnFlow and DiffKG, KG paths are predicted by scoring entity paths and relation paths respectively, and then rerank which makes it harder to achieve optimum. While a KG triple is composed of both, our model uses PA-MLP to aggregate all the exit path information to improve the perception of the agent, which is a more reasonable modeling approach.
HiTKG is state-of-the-art KG walker, which build a multi-hierarchy attention block to aggregate the multiscale information. However, different types of input data sources are difficult to aggregate, and how well they are aggregated directly affects the performance of the model. We unify all the multi-scale input sources into the prompt, and due to the large model has a large number of instruction comprehension ability to get a rich information encoding representation.
The GPT-4-Standard and GPT-4-Normal methods are deficient in path awareness. GPT-4-OPA exhibit improved outcomes with the addition of path awareness. The generation of GPT-4 paths is entirely dependent on the background knowledge in the dataset during the training phase, and the GPT-4 generative model itself is not designed for sequential decision-making tasks, and achieving such a score has impressed us. Although we added states to GPT-4 through prompt, this is limited by the length of the prompt, which is fundamentally due to the fact that GPT-4 is inherently memoryless. Based on these factors, optimizing GPT-4 on multi-hop inference datasets of the Knowledge Graph to further improve its performance, generating human-preferred inference paths on large-scale Knowledge Graph datasets is still a challenge.
Considering LLMs as agents that explore a knowledge graph to acquire experience can benefit benefits from the positive-negative feedback optimization mechanism of the Reinforcement Learning Policy Supervisor Algorithm. This method enhances the training of our model to perform flexible reasoning on KGs in multi-step scenarios, outperforming not only GPT-4 but also smaller models. The model’s superior performance corroborates the effectiveness of our approach.
@@
LLaMA-2-7B-ARK 16.59 27.17 34.85 47.88 62.32 36.39 53.63 65.68 80.20 89.68
LLaMA-2-7B-ARK-UI 16.87 27.01 34.35 47.67 63.03 34.70 52.02 62.66 78.09 88.65
LLaMA-2-7B-ARK-FP32 14.59 24.64 32.24 45.80 61.75 34.35 52.57 62.51 79.18 88.51
LLaMA-2-7B-ARK-WT 1.10 3.44 5.71 10.47 15.42 9.45 19.46 50.94 71.43 94.03
LLaMA-2-7B-ARK-WP 0.98 2.53 3.28 4.88 5.52 18.90 40.67 54.76 77.25 93.08
@@
4.6 Analysis Experiment
As shown in Table 2, LLM-ARK was benchmarked against multiple ablation models on the OpenDialKG dataset. (1) First, to evaluate the impact of instructions on model performance, we trained the LLaMA-2-7B-ARK-UI model without instruction. The results of this model are the closest to those of LLaMA-2-7B-ARK, indicating that the presence or absence of commands has an effect on the model’s results, but not a serious one. (2) Then, we have implemented IEEE 754 floating-point format (FP32) operations for our experiments. The results show that using the BFLOAT16 tensor for training, recall@K gives better results than FP32 without changing the hyperparameters. (3) Next, LLaMA-2-7B-ARK-WT that was trained by randomly initializing relation and entity embedding. The decrease in performance indicates that the absence of knowledge graph representation learning has negatively impacted the training process. (4) We conducted the fourth ablation experiment LLaMA-2-7B-ARK-WP and found that the performance of our export environment-aware sub-module PA-MLP decreases substantially if we do not consider the exit paths. The score for path@1 is only 0.98%, which is 15.61% lower than LLaMA-2-7B-ARK. These ablation experiments and results demonstrate the contribution and necessity of our subcomponents to the model.
4.7 Case Study
@@ 5 Conclusion
This paper evaluates the ability of current state-of-the-art LLM-based dialog systems in handling KG conversational reasoning tasks. To enhance LLM’s performance on this task, we introduce LLM-ARK, a full-text environment-aware Knowledge Graph inference agent optimized using online reinforcement learning. Empirical analysis demonstrates that our model outperforms GPT-4 and smaller models. The experiments also sheds light on the model’s performance can be seriously affected by the mismatch between the LLMs and the environment information. Our method can inspire subsequent researchers to pay attention to the critical role of considering various factors during model optimization in the field of LLM-based conversational KG reasoning tasks.
@@ A.2 Tricks
In the original paper of PPO, no implementation details and techniques are mentioned other than the use of GAE to compute the dominance function. Referring to this repository [^3] https://github.com/Lizhi-sjtu/DRL-code-pytorch, we employ several optimization tricks. In the actual code implementation, to encourage the diversity of paths sampled by the strategy during training, we added an entropy regularization term to our loss function. We used the operation of normalization of advantage proposed in the paper [Tucker et al., 2018]. Learning rate decay can enhance the smoothness in the late training stage to some extent and improve the training effect. Here we use the linear decay of learning rate, with the number of training steps learning rate from the initial value of a linear decline to 0. Gradient clipping is a trick introduced to prevent the gradient from exploding during the training process, which also serves to stabilize the training process. Orthogonal Initialization is a neural network initialization method proposed to prevent problems such as gradient vanishing and gradient explosion at the beginning of training. Referring to the MAPPO [Yu et al., 2022], the Adam optimizer individually sets eps=1e-5, and this particular setting can improve the training performance of the algorithm to some extent.
A.3 Limitation
A.3.1 Limitations of Inference Efficiency
Efficiency is always a significant issue when building deep learning models based on LLMs. Although our research freezes the parameters of the LLM in the back-propagation stage and uniformly uses the bfloat16 computational type, the huge number of parameters of the model leads to inefficient forward propagation and large GPU memory usage when collecting experience and inference. As stated in the LLaMA paper[Touvron et al., 2023a], the efficiency of the model’s inference is more crucial than its training efficiency. It is acceptable for the training process to be slower, but the inference must be faster. Improving the inference speed of the model while ensuring its effectiveness is a challenge. In addition to constructing the research model, online applications based on LLMs must also address the efficiency issue. Conversational reasoning models based on LLMs must be efficient for real-time applications. The inference efficiency is crucial for building online applications based on LLMs.
A.3.2 Limitations of Entity Embedding
Our research work has identified limitations in the semantic representation of knowledge graph entities. The attributes of knowledge graph entities should be considered during the reasoning process. However, these attributes may be lengthy descriptions that are not easily processed by our TransE knowledge graph semantic embedding model. Furthermore, while most knowledge graphs are currently represented in text form, it is equally important to consider multimodal knowledge graph reasoning in research. By constructing a reasoning model based on multimodal inputs, machines can better describe and understand the real world. For example, the soon-to-be two-dimensionalized Law in Three Body Death Forever says ”Oh, it’s time to go into the picture, kids, go ahead,” and the user asks a question about this scenario, ”can you help me find some pictures related to this galaxy?”. The model may need to deduce that the two-dimensional representation portrayed in this book is the Milky Way galaxy, and then locate relevant images of the galaxy. Our model is currently unable to incorporate the combination of multi-modal, multi-attribute entities, which is a limitation of our work in this endeavor, as well as an area for future research efforts.
A.4 GPT4 Prompts
For GPT4-OPA prompt, since max length is 2, we need to recursively get all exit paths at the next level of all exit paths of the current entity, most of which are omitted due to the large number of KG subgraph triples of exit paths.
@@ A.4 GPT4 Prompts
### Environment
Current Entity: Washington Redskins
@@
Path History: []
Current Entity: Washington Redskins
@@
Path History: []
Current Entity: Washington Redskins
@@ A.5 HyperParameters
number of explorations 8
replay buffer size 4096